OpenStudy (anonymous):

Are 24, 45, and 51 the side measurements of a special right triangle? Yes or No, Explain your answer and show work.

OpenStudy (anonymous):

you dont have to show work just need to know why ..

OpenStudy (cwrw238):

check if 51^2 = 45^2 + 24^2

OpenStudy (cwrw238):

or more simply 17^2 = 15^2 + 8^2 its all about the pythagoras theorem

OpenStudy (anonymous):

2601=2025+576 2601=2601

OpenStudy (cwrw238):

that proves its right angled

OpenStudy (anonymous):

but is it a special right triangle ?

OpenStudy (cwrw238):

these numbers are called pythagorean triads - yes special because the sides are all whole numbers the simplest one is 3-4-5